OpenStudy (anonymous):

PLEASEEE HELP!!! What is the domain of (√x+8)/(x+4)(x-6) x ≠ -8, x ≠ -4, x ≠ 6 x > 0 All real numbers x ≥ -8, x ≠ -4, x ≠ 6

OpenStudy (anonymous):

@KlOwNlOvE @tkhunny @SolomonZelman

OpenStudy (anonymous):

square roots need to have values which are >= 0 and denominator cannot have any vaule which make them 0

OpenStudy (anonymous):

so x+8>=0 and x is not -4 and x is not 6

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

I love you a lot! lol thank you

OpenStudy (anonymous):

Do you think you can help me with one more?

OpenStudy (anonymous):

ur welcome:) sure, if i can

OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (anonymous):

(x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150 f(x) = x4 - 9x2 - 50x - 150

OpenStudy (anonymous):

do you know about complex numbers ?

OpenStudy (anonymous):

they are of the form a + bi, and i^2=-1 you should know it

OpenStudy (anonymous):

and it is a rule that if a polynomial has a 0 which is a+bi, then the other 0 should be a-bi, then here we have -1+3i so the other 0 should be -1+3i, and others are 5 and -3 as given

OpenStudy (anonymous):

lets add them: -5 + (-3) + (-1+3i) + (-1-3i) = -10

OpenStudy (anonymous):

so the sum of 0s are -10

OpenStudy (anonymous):

What do I do with the -10?

OpenStudy (anonymous):

yeah i was looking into it, ok so we can construct an equation in the form (x+a)(x+b)(x+c)(x+d)

OpenStudy (anonymous):

lets try it: we have all the zeros so it will be like: (x+5)(x+3)(x+1-3i)(x+1+3i)

OpenStudy (anonymous):

we should now multiply them.. oops by the way the form is (x-a)(x-b)(x-c)(x-d)=0 if the 0s are a,b,c and d

OpenStudy (anonymous):

so lets first multiply the first two terms: (x+5)(x+3)=x^2+8x+15

OpenStudy (anonymous):

Wouldn't they be negative though?

OpenStudy (anonymous):

then lets multiplty the last two: ((x+1)-3i)((x+1)+3i) = ((x+1)^2)+3i(x+1) - 3i(x+1) -9i^2

OpenStudy (anonymous):

= x^2+1+2x + 9

OpenStudy (anonymous):

so we can now multiply the first result with this: (x^2+8x+15)(x^2+2x+10) = x^4+2x^3+10x^2 + 8x^3+16x^2+80x + 15x^2+30x+30 = x^4+10x^3+26x^2+110x+30

OpenStudy (anonymous):

sth is wrong...

OpenStudy (anonymous):

yeaaaa it should be x-5 above, i worte x+5

OpenStudy (anonymous):

That isn't one of my options though

OpenStudy (anonymous):

x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150 f(x) = x4 - 9x2 - 50x - 150

OpenStudy (anonymous):

let me revise: (x-5)(x+3) = x^2-2x-15

OpenStudy (anonymous):

so (x^2-2x-15)(x^2+2x+10) = (x^4+2x^3+10x^2 - 2x^3 -4x^2-20x - 15x^2-30x-150 = x^4 -9x^2-50x-150

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Thank you for all your help. Youre the best :))

OpenStudy (anonymous):

np, but please notice what we did: we first thought that if a polynomial has a complex 0 in it in the form a+bi, then the other 0 should be a-bi; and if we multiply all the 0s then we can get the equation.. know that the equation is of the form (x-a)(x-b)(x-c)(x-d)=0 where a, b, c, and d are all 0s (the 3 of which are given in the question and the last one we found as -1-3i) and we multiplied them all.. Notice how i multiplied the complex numbers: i grouped the real parts and imaginary parts together and multiplied them that way and also know that i^2 = -1

OpenStudy (anonymous):

I understand :) @barbillus